OpenStudy (anonymous):

x^5 + x^3 + 8x^2 + 8 factor

OpenStudy (anonymous):

so far i got [x^3(x^2+1)+ 8(x^2+1)] (x^2+1)(x^3+8)

OpenStudy (jackellyn):

Well theres a formula for (x^3+8) a3 + b3 = (a + b)(a2 – ab + b2) a3 – b3 = (a – b)(a2 + ab + b2) Would yours be the top one or the bottom one?

OpenStudy (anonymous):

the bottom

OpenStudy (jackellyn):

It would be the top one because you are adding, right?

OpenStudy (anonymous):

your right

OpenStudy (jackellyn):

so what is the cube root of 8?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

(x-2)(x2+2x+4)

OpenStudy (jackellyn):

remember you're using the first formula\ so it should be (a+b)(a^2-ab+b^2)

OpenStudy (jackellyn):

you have all the numbers right but the signs need to be changed

OpenStudy (anonymous):

(x+2)(x2-2x+4)(x2+1)

OpenStudy (jackellyn):

right! no what about (x^2 +1) can you factor that?

OpenStudy (anonymous):

x= +_ i .. no?

OpenStudy (jackellyn):

right

OpenStudy (jackellyn):

so (x+i)(x-i)

OpenStudy (jackellyn):

so what's the whole thing factored out?

OpenStudy (anonymous):

(x+2)(x2-2x+4)(x+i)(x-i)

OpenStudy (jackellyn):

I have one more question from you, can you factor (x^2-2x+4)?

OpenStudy (anonymous):

no would you use the quadratic formula

OpenStudy (jackellyn):

very good!

OpenStudy (anonymous):

(x+2)(x+(-1+i sqrt{3})(x+(-1-i sqrt{3})(x+i)(x-i)

OpenStudy (jackellyn):

yeah, thats about right.

OpenStudy (anonymous):

thx for your help jackellyn!!!

OpenStudy (jackellyn):

no problem :)